Given the cost function C(x) and the revenue function R(x)​, find the number of units x that must be sold to break even.
C(x)=13x+80,000 and R(x)=18x.

Answer by ikleyn(52781)

.
Given the cost function C(x) and the revenue function R(x)​, find the number of units x
that must be sold to break even.
C(x)=13x+80,000 and R(x)=18x.
~~~~~~~~~~~~~~

The break even, by the definition, is when Cost = Revenue,  or

    13x + 80000 = 18x.


From this equation

    80000 = 18x - 13x

    80000 =    5x

        x =    80000/5 = 16000.


ANSWER.  16000 units should be sold.

Solved.